In a binomial distribution with an n of 15 and a probability of success of 20 percent, what is the probability of 0? 9) OA 0.3520 OB 0.0352 ОС 0.1319 OD 02309 The travel times of trains between station A and station B are uniformly distributed with a low of 15 minutes and a high of 21 minutes. What is the probability of the travel time being 16 minutes or more? (10) OA 0.333 OB 0.167 ОС 0.833 OD 0.154 A coffee vending machine is programmed to pour 8 ounces of coffee into a cup. According to the manufacturers specifications the population standard deviation of each pour is 0.3 ounces, and the distribution is normal. What is the probability of randomly getting a cup of coffee and finding it contains between 8.3 and 8.6 ounces? (11) OA 0.3413 OB 0.4772 С 0.81 85 OD 0.1359 (12) The travel times of trains between station A and station B are uniformly distributed with a low of 15 minutes and a high of 21 minutes. What is the probability of the travel time being 20 minutes or more? OA 0333 OB 0.167 ОС 0.833 OD 0.154
Answer #1

9) n = 15

p = 0.2

P(X = 0) = 15C0 * 0.2 * (1 - 0.2)15-0 = 0.0352

Option-B) 0.0352

10) X ~ U (15 , 21)

P(X > 16) = (21 - 16) / (21 - 15) = 0.833

Option-C) 0.833

11) µ = 8

sd = 0.3

P(8.3< X< 8.6) P(

                                        =P(\frac{8.3-8}{0.3}<Z<\frac{8.6-8}{0.3})

                                        = P(1 < Z < 2)

                                        = P(Z < 2) - P(Z < 1)

                                        = 0.9772 - 0.8413

                                        = 0.1359

Option-D) 0.1359

12) X ~ U(15 , 21)

P(X > 20) = (21 - 20) / (21 - 15) = 0.167

Option-B) 0.167